Find the eigenvalues and corresponding eigenvectors of matrix G given below.

G= [2 -3]

[4 -5]


1
Expert's answer
2021-09-10T14:35:02-0400

"G= \begin{bmatrix}
2 & -3 \\
4 & -5
\end{bmatrix}"

Finding the eigenvalues:

"\det \left(\begin{pmatrix}2&-3\\ 4&-5\end{pmatrix}-λ\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)"

"=\begin{pmatrix}2-λ&\left(-3\right)-0\\ 4-0&\left(-5\right)-λ\end{pmatrix}"

"\left(2-λ\right)\left(-5-λ\right)-\left(-3\right)\cdot \:4"

"\u03bb^2+3\u03bb+2=0"

Solving this "\u03bb^2+3\u03bb+2=0"

"λ=-1,\:λ=-2"

Therefore eigenvalues are: -1 and -2

Calculating Eigenvectors for ": \u03bb=-1"

"= \begin{pmatrix}2&-3\\ 4&-5\end{pmatrix}-\left(-1\right)\begin{pmatrix}1&0\\ 0&1\end{pmatrix}"

"=\begin{pmatrix}2&-3\\ 4&-5\end{pmatrix}-\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}"

"=\begin{pmatrix}3&-3\\ 4&-4\end{pmatrix}"

To solve "\begin{pmatrix}3&-3\\ 4&-4\end{pmatrix} \begin{pmatrix} x\\y\end{pmatrix} = \begin{pmatrix} 0\\0\end{pmatrix}," reduce the matrix

"= \begin{pmatrix} 1\\1 \end{pmatrix}" as the eigenvectors for "\u03bb=-1"

Calculating Eigenvectors for ": \u03bb=-2"

"= \begin{pmatrix}2&-3\\ 4&-5\end{pmatrix}-\left(-2\right)\begin{pmatrix}1&0\\ 0&1\end{pmatrix}"

"=\begin{pmatrix}4&-3\\ 4&-3\end{pmatrix}"

To solve "\begin{pmatrix}4&-3\\ 4&-3\end{pmatrix} \begin{pmatrix} x\\y\end{pmatrix} = \begin{pmatrix} 0\\0\end{pmatrix}," reduce the matrix

"= \begin{pmatrix} 3\\4 \end{pmatrix}" as the eigenvectors for "\u03bb=-2"

The Eigenvectors for "\begin{pmatrix}2&-3 \\ 4 & -5\end{pmatrix}" becomes "=\begin{pmatrix}1\\ 1\end{pmatrix},\:\begin{pmatrix}3\\ 4\end{pmatrix}"